Three consecutive numbers of G.p are such that there sum is 26 and there product is 216. Find the numbers.
Answer by richwmiller(17219)

a = t * r^(n-1),
b = t * r^(n),
c = t * r^(n+1),
a+b+c=26,
a*b*c=216
a=2, b=6, c=18, r=3, t1=6
or
a=18, b=6, c=2, r=1/3, t1=18